OpenStudy (linn99123):

Use an addition property to solve for a. 10 + (–15) = –15 + a A. –25 B. –10 C. –5 D. 10

OpenStudy (lieutenantgeneral):

10

OpenStudy (linn99123):

awwww thank you sweet face

OpenStudy (abb0t):

10 - 15 = -15 + a. All you do is remove the parenthesis and think of it as subtraction. 10-15 = -5 so you have: -5 = -15 + a now, add 15 to both sides, and you should get a = 10 vote for me: http://www.easypolls.net/poll.html?p=5421db85e4b09149347e0ab3

OpenStudy (linn99123):

Ight thanks and abo0t add me to that easypolls thing PELASE